OpenStudy (anonymous):

What is the first step when solving 4x + 10 = 2y for x?

OpenStudy (anonymous):

Putting y as 0, if you are finding the x-intercept..

OpenStudy (anonymous):

So Basically Elimination: 1st: 2x+3y=-16 2nd: 5x-10y=30 Multiply 1st by 5 and 2nd by 2 to get the following: 1st: 10x+15y=-80 2nd: 10x-20y=60 Now subtract the 2nd from the 1st to eliminate the x's: 35y = -140 y = -4 Substitute that value of "y" into the original 1st equation to solve for "x": 2x+3(-4)= -16 2x-12 = -16 x = -2 Check this answer of x=-2 and y=-4 in the 2nd original equation: 5(-2)-10(-4)= 30 -10+40 = 30 30 = 30

OpenStudy (anonymous):

@babygirlabi start by subtracting 10 from both sides, then divide by 4

OpenStudy (anonymous):

thxxxx

OpenStudy (anonymous):

-_-

OpenStudy (yellowlegoguy99):

lel you copied and pasted that stuff, ldskbgkhbcd

OpenStudy (anonymous):

the question was to isolate x no find the x-intercept

OpenStudy (anonymous):

not*